When i cycled from home at a speed at 36km/h i reached the station 4mins after the train left.if i had cycled at 45km/h i would reach to the station 4min before the train left.how far is the station
1
Expert's answer
2020-04-23T12:58:55-0400

Let x be the time from leaving the house to the departure of the train (in hours); (4 / 60) - converting minutes to hours. Then 36 × (x + (4 / 60)) is a distance from home to the station; 45 × (x - (4 / 60)) is also distance from home to the station, so we can write

36 × (x + (4 / 60)) = 45 × (x - (4 / 60))

Solving this equotion we get

x = 0.6 h = 36 (min)

And so the distance is

36 × (0.6 + (4 / 60)) = 24 (km)